OpenStudy (anonymous):

Solve. x(2) + x – 6 = 0

OpenStudy (anonymous):

Is the first X squared?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

X(2) +X-6=0 X(2) + X-6+6= 0+6 x(2) + X = 6 2X(2)=6 2X(2) / 2 = 6/2 X(2)= 3 X = sqrt 3 It could be this